OpenStudy (anonymous):

How does 1^(-n) * (-1)^n simplify to (-1)^n?

OpenStudy (campbell_st):

well if you look at the 1st term \[1^{-n} = \frac{1}{1^n}\] this is always positive and is just 1 and (-1)^n alternates between 1 and -1 so it becomes \[1 \times (-1)^n\] which is \[1 \times (-1)^n = (-1)^n \]